if a cube's length is 4 1/2 ft. what is the surface area of the cube?
Answer by rosey51235(30)

Hi:)
First, you need the surface area of a single face of the cube, so first put 4 1/2 into an improper fraction to be multiplied. It is 9/2, so 9/2*9/2=81/4, whicc reduces to 20 1/4, so the area of one face of the cube is 20 1/4 square feet.
To find the total surface area, there are 6 faces on a cube, so just multiply the single face area by 6. (must be improper fraction to multiply) 81/4*6/1=486/4, or 121 1/2.
But, most times an answer is expected as a decimal, so the total surface area of the cube is 121.5 square feet.
Hope I helped!:)
-Rose